OpenStudy (anonymous):

Factor the fallowing Help plz :( Y=4x2-32x+28 Y=x^2-36 Y=x^2-2x-3

OpenStudy (anonymous):

all 3?

OpenStudy (anonymous):

Yes Please (:

OpenStudy (anonymous):

Y = 4 (x-7) (x-1) Y = (x-6) (x+6) Y = (x-3) (x+1)

OpenStudy (anonymous):

UGH TY I KNOW HOW TO DO IT (: know (:

OpenStudy (anonymous):

ok, np

OpenStudy (anonymous):

um they are rong answers?

OpenStudy (anonymous):

Wait sorry wrng person..

OpenStudy (anonymous):

oh, lol